# Thompson Sampling-Based Learning and Control for Unknown Dynamic Systems

### v1

@@ IV-A Parameterization of Controllers
In the following, we show that the constructed function space can be quantitatively evaluated by analyzing the regret of the optimal control law $\bar{\bm{g}}=\arg\min\limits_{{\bm{g}}\in\mathcal{G}}J$ with respect to to the optimal control law ${\bm{g}}^{*}$ in Assumption 3.
@@
A reproducing kernel Hilbert space (RKHS) $\mathcal{H}(K)$ is a Hilbert space of functions on a set $\mathcal{X}$ with a reproducing kernel $K(\cdot,\cdot)$ that satisfies the following properties:
@@
    $K({\bm{x}},\cdot)\in\mathcal{H}(K)$ for all ${\bm{x}}\in\mathcal{X}$;
@@
    $\langle g,K({\bm{x}},\cdot)\rangle_{\mathcal{H}(K)}=g({\bm{x}})$ for all ${\bm{x}}\in\mathcal{X}$ and $g\in\mathcal{H}(K)$,
[…]
@@
Let Assumptions 1 and 3 hold. Under the proposed parameterization approach (12), the constrcted learning set $\mathcal{G}$ satisifes: $\exists\bar{\bm{g}}=\arg\min\limits_{{\bm{g}}\in\mathcal{G}}J$:
@@
Specifically, according to the definition of $\mathcal{G}_{i}$ in 4, it can be verified that for $\forall{\bm{w}}\subseteq[1,n]$, the function $g^{(i)}_{{\bm{w}},\tilde{\bm{x}}}$ is in the RKHS $\mathcal{H}(K_{n,{\bm{w}}}^{(i)})$ with the reproducing kernel $K_{n,{\bm{w}}}^{(i)}({\bm{x}},\tilde{\bm{x}})$. Here, $\mathcal{H}(K_{n,{\bm{w}}}^{(i)})$ is a Hilbert space of real functions defined on $\mathcal{X}_{\bm{w}}:=\{{\bm{x}}_{\bm{w}}|x\in\mathcal{X}\}$.
According to the definition (IV-A), functions $g\in\mathcal{G}_{i}$ can be regarded as linear combinations of base functions $g^{(i)}_{{\bm{w}},\tilde{\bm{x}}},~{}{\bm{w}}\subseteq[1,n]$. Therefore, for $g\in\mathcal{G}_{i}$, a RKHS can be obtained as $\mathcal{H}(K_{n}^{(i)})$ with the reproducing kernel $K_{n}^{(i)}({\bm{x}},\tilde{\bm{x}})$ being
@@
Theorem 1 analyzes the regret introduced by the proposed controller parameterization method. Assumption 3 assumes an optimal control law ${\bm{g}}^{*}$ in the space $\mathcal{G}^{*}$. Meanwhile, the parameterization method (12) constructs a convex hull $\mathcal{G}_{i}$. The reduction of the feasible space inevitably leads to changes in the optimal solution of the optimization problem, which is quantified in Theorem 1 through the regret $\|{\bm{g}}^{*}-\bar{\bm{g}}_{n}\|_{\mathcal{L}_{2}}$. Theorem 1 demonstrates that the proposed controller parameterization method ensures the boundedness of the regret.
Remark 5.
The regret of the parameterization method is guaranteed by the properties of the Hilbert space (Lemma 1). Therefore, the proof of Theorem 1 primarily aims to demonstrate that the constructed function space is a Hilbert space. The proposed approach not only ensures the boundedness of the learning regret through Lemma 1 but also provides an advantageous structural foundation for solving optimization problems within the function space.
Using Theorem 1, the error bound for vector-valued functions can be further derived, which is presented in the following corollary.
@@
Let Assumptions 1 and 3 hold. Under the proposed parameterization approach (12), the constructed learning set $\mathcal{G}$ satisfies: $\exists\bar{\bm{g}}=\arg\min\limits_{{\bm{g}}\in\mathcal{G}}J$:
$\|\bar{\bm{g}}-{\bm{g}}^{*}\|^{2}_{\mathcal{L}_{2}}\leq\frac{mM_{g}^{2}-\sum_{i=1}^{m}\|{g}_{i}^{*}\|_{\mathcal{L}_{2}}^{2}}{2^{n}}.$ (18)
@@ IV-B TS-based Learning Algorithm
For the considered system (1) and a given data segment $\{{\bm{x}}_{t}(0),{\bm{u}}_{t}(0),\ldots,{\bm{x}}_{t}(K),{\bm{u}}_{t}(K)\}$, the control cost can be calculated according to its definition. However, due to the unknown system dynamics $f(\cdot)$ and uncertain noise ${\bm{w}}$, the cost function $J$ is not available unless the control law is executed. This brings challenges to the controller design. To address this issue, we consider a function space $\bar{\mathcal{J}}$ with $\sigma$-field $\Sigma_{\bar{\mathcal{J}}}$ and measure $v_{\bar{\mathcal{J}}}(\cdot)$ such that the function $\bar{J}({\bm{g}};{\bm{x}})\in\bar{\mathcal{J}}$ satisfies the following properties:
@@
    Normalization:
    $\int_{\mathcal{G}}\bar{J}({\bm{g}};{\bm{x}}){\rm d}v_{\bar{\mathcal{J}}}({\bm{g}})=1;$ (20)
@@
    $\mathbb{P}({\bm{g}}\in\Omega)=\int_{\Omega}\bar{J}({\bm{g}};{\bm{x}}){\rm d}v_{\bar{\mathcal{J}}}({\bm{g}}),~{}\forall\Omega\subseteq\mathcal{G};$ (21)
@@
    $\bar{J}({\bm{g}};{\bm{x}})\propto\frac{1}{J},~{}{\bm{g}}\in\mathcal{G},{\bm{u}}(k)={\bm{g}}({\bm{x}}(k)),{\bm{x}}(0)={\bm{x}}.$ (22)
@@
    Boundedness: $v_{\bar{\mathcal{J}}}(\bar{\mathcal{J}})=\bar{M}_{\bar{\mathcal{J}}},~{}\underline{J_{b}}\leq\bar{J}({\bm{g}};{\bm{x}})\leq\overline{J_{b}}.$
Here we note that, through the aforementioned five properties, we establish a space $\bar{\mathcal{J}}$, which serves as a tool for indirectly learning the unknown cost function $J(\cdot)$. Although there are various possibilities for the type of the unknown cost function, under the boundedness assumption 2, we can construct a function $\bar{J}({\bm{g}};{\bm{x}})$ belonging to the space $\bar{\mathcal{J}}$ and related to the cost function $J(\cdot)$. This facilitates the following learning mechanism design and learning convergence analysis.
@@
$\bar{J}^{t}:=\mathbb{E}\left[\bar{J}\right]:=\int_{\bar{\mathcal{J}}}\bar{J}F_{\rm pdf}^{t}({\rm d}v_{\mathcal{J}}(\bar{J})).$ (27)
@@
Using the proposed distribution function $F_{\rm pdf}^{t}(\bar{J})$, the TS based learning algorithm can be designed and is summarized in Algorithm 2. After obtaining the initial data (Lines 1-5), the control law is updated at the beginning of each segment. Specifically, at the beginning of each segment, the control cost $J_{t-1}({\bm{g}}_{t-1})$ is calculated using the data collected in the previous segment, and $R_{t}$ and the posterior distribution function $F_{\rm pdf}^{t}(\bar{J})$ are updated (Line 7). Here, a small trick is employed: although $\bar{J}^{*}({\bm{g}}_{t})$ and $1/J_{t}({\bm{g}}_{t})$ differ by a proportional coefficient (as shown in (25)), this coefficient appears in both the numerator and denominator when updating the posterior distribution, so it does not affect the calculation result. Then, a function $\bar{J}_{t}$ is randomly sampled according to the posterior distribution function $F_{\rm pdf}^{t}(\bar{J})$ (Line 8), and the control law ${\bm{g}}_{t}$ is selected to maximize $\bar{J}_{t}({\bm{g}}_{t})$ (Line 9). Finally, the control law ${\bm{g}}_{t}$ is applied, and the system data for the $t$-th segment $\{{\bm{x}}_{t}(0),{\bm{u}}_{t}(0),\ldots,{\bm{x}}_{t}(K),{\bm{u}}_{t}(K)\}$ is collected (Line 10).
@@ IV-C Convergence Analysis of the Learning Process
$T_{d}(x)$ :=$:=\left\{\!\!\begin{array}[]{ll}\log x,&\text{if}~{}d=D_{\rm KL},\\ \sqrt{x}-1,&\text{if}~{}d=\frac{1}{2}D^{2}_{\rm H}.\end{array}\right.$ (31)
@@
[…]
@@
$\lim\limits_{t\rightarrow\infty}\inf d(\bar{J}^{t},\bar{J}^{*})>0;$ (34)
@@
The proof of this theorem is divided into two steps. In the first step, a martingale process is constructed, and several properties of the martingle are analyzed. In the second step, the proof of Theorem 2 is completed by using the properties of the constructed martingale process. In the following analysis, the complementary set of the neighborhood ${\mathcal{B}}(\bar{J}^{*},\delta_{\bar{J}})$ is denoted as $\Omega={\mathcal{B}}^{c}(\bar{J}^{*},\delta_{\bar{J}})$.
@@
[…]
@@
[…]
@@
$\log~{}L_{T}(\Omega)-\log~{}L_{0}(\Omega)$ $<-T\epsilon_{L}$ (53)
$L_{T}(\Omega)$ $<L_{0}(\Omega)\exp(-T\epsilon_{L}),$ (54)
@@
[…]
@@
Remark 6.
Theorem 2 characterizes the convergence of the predictive density $\bar{J}^{T}$ under the proposed learning algorithm. Specifically, as time progresses, $\bar{J}^{T}$ converges to a small neighborhood ${\mathcal{B}}(\bar{J}^{*},\delta_{\bar{J}})$. In the proof of Theorem 2, the convergence result is established by constructing a martingale process and applying the martingale convergence theorem to show that the probability of $\bar{J}^{T}$ belonging to the complement of the neighborhood of $\bar{J}^{*}$ decreases exponentially. This neighborhood is defined by $\bar{J}^{*}$ and a small positive constant $\delta_{\bar{J}}$, where $\bar{J}^{*}$ depends on the system dynamics and noises (as shown in (25)). The parameter $\delta_{\bar{J}}$ is a small positive constant that characterizes the size of the neighborhood. As indicated in (51) and (56), the parameter $\delta_{\bar{J}}$ affects the convergence rate $\epsilon_{L}$, as $\delta_{\bar{J}}$ determines the set $\Omega={\mathcal{B}}^{\rm c}(\bar{J}^{*},\delta_{\bar{J}})$.
Remark 7.
The convergence of Theorem 2 holds for both the Hellinger distance and the KL divergence, which are two widely used methods for describing the similarity between different pdfs. On the one hand, discussing both metrics is helpful because the convergence of one metric does not necessarily guarantee the convergence of the other. On the other hand, the proof approach of the proposed method has the potential to be extended to other metrics, such as the Wasserstein distance, Jensen-Shannon divergence, and Bhattacharyya distance, etc. For these metrics, the core of the proof would lie in constructing an appropriate function $T_{d}(\cdot)$ and the corresponding martingale process $M_{T}$, and then deriving the convergence of the learned function $\bar{J}^{t}$ through the convergence of the martingale process.

### v2

@@ IV-A Parameterization of Controllers
We recall that $\bm{g}$$^{*}$ is the optimal control law, and $\bar{\bm{g}}$ is the approximated optimal control law within the function space $\mathcal{G}$ which satisfies $\bar{\bm{g}}=\arg\min\limits_{{\bm{g}}\in\mathcal{G}}J$. The constructed function space $\mathcal{G}$ can be quantitatively evaluated by analyzing the regret of $\bar{\bm{g}}$ with respect to the optimal control law $\bm{g}$$^{*}$, which is analyzed as follows.
@@
A reproducing kernel Hilbert space (RKHS) $\mathcal{H}({\rm Ker})$ is a Hilbert space of functions on a set $\mathcal{X}$ with a reproducing kernel ${\rm Ker}(\cdot,\cdot)$ that satisfies the following properties:
@@
    ${\rm Ker}({\bm{x}},\cdot)\in\mathcal{H}({\rm Ker})$ for all ${\bm{x}}\in\mathcal{X}$;
@@
    $\langle g,{\rm Ker}({\bm{x}},\cdot)\rangle_{\mathcal{H}({\rm Ker})}=g({\bm{x}})$ for all ${\bm{x}}\in\mathcal{X}$ and $g\in\mathcal{H}({\rm Ker})$,
where $\langle\cdot,\cdot\rangle$ represents the inner product in $\mathcal{H}({\rm Ker})$.
@@
Let Assumptions 1 and 3 hold. Under the proposed parameterization approach (15), the constructed learning set $\mathcal{G}_{i}$ satisfies: $\exists\bar{\bm{g}}=\arg\min\limits_{{\bm{g}}\in\mathcal{G}_{i}}J$:
@@
[…]
@@
Theorem 1 is an existence result. Specifically, in this work, we construct a function space $\mathcal{G}$ and aim to find a control law $\bar{\bm{g}}$ in $\mathcal{G}$ to approximate the optimal control law $\bm{g}$$^{*}$. By analyzing the properties of the function space $\mathcal{G}$, Theorem 1 proves the existence of $\bar{\bm{g}}$ and the boundedness of regret. The specific method for finding the control law $\bar{\bm{g}}$ in $\mathcal{G}$ will be described in detail in subsequent sections of this work.
Theorem 1 addresses the conservatism in approximating the optimal control law within the constructed function space $\mathcal{G}$. This conservatism can be considered from two aspects. First, the construction of $\mathcal{G}$ is determined by the function space spanned by the initial control law, rather than the specific law itself. Thus, even if the initial control law is unstable, as long as the spanned space is rich enough, it is still possible to learn a high-quality control law within $\mathcal{G}$. Second, the conservatism is quantitatively characterized by the distance between $\mathcal{G}$ and the optimal control law $\bm{g}$$^{*}$. Specifically, we introduce the reachable error $\Delta_{\mathcal{G}}:=\inf_{{\bm{g}}\in\mathcal{G}}\|{\bm{g}}-{\bm{g}}^{*}\|_{\mathcal{L}_{2}}$ to capture the approximation error induced by the limitations of the chosen function space. Theorem 1 shows that this reachable error is finite under the proposed construction. Moreover, by increasing the dimension of $\mathcal{G}$, the upper bound of the reachable error can be reduced, thereby alleviating the conservatism introduced by the initial function space selection. The reachable error $\Delta_{\mathcal{G}}$ is provided in the following corollary.
@@
Let Assumptions 1 and 3 hold. Under the proposed parameterization approach (15), the reachable error of the constructed function space $\mathcal{G}$ satisfies
$\Delta_{\mathcal{G}}\leq\sqrt{\frac{mM_{g}^{2}-\sum_{i=1}^{m}\|{g}_{i}^{*}\|_{\mathcal{L}_{2}}^{2}}{2^{n}}}.$ (20)
@@ IV-B TS-based Learning Algorithm
For the considered system (1) and a given data segment $\{{\bm{x}}_{t}(0),{\bm{u}}_{t}(0),\ldots,{\bm{x}}_{t}(K_{t}),{\bm{u}}_{t}(K_{t})\}$, the control cost can be calculated according to its definition (namely, (2) or (3)). However, due to the unknown system dynamics $f(\cdot)$ and uncertain noise $\bm{w}$}, the cost function $J$ is not available unless the control law is executed. This brings challenges to the controller design. To address this issue, we consider a function space $\bar{\mathcal{J}}$ with $\sigma$-field $\Sigma_{\bar{\mathcal{J}}}$ and measure $v_{\bar{\mathcal{J}}}(\cdot)$ such that the function $\bar{J}({\bm{g}};{\bm{x}}_{0})\in\bar{\mathcal{J}}$ satisfies the following properties:
@@
    Probability measure:
    $\int_{\mathcal{G}}\int_{\mathcal{X}}\bar{J}({\bm{g}};{\bm{x}}_{0}){\rm d}v_{\bar{\mathcal{J}}}({\bm{g}}){\rm d}{\bm{x}}_{0}=1;$ (22)
@@
    $\forall\Omega\subseteq\mathcal{G},~\forall\Xi\subseteq\mathcal{X}:$
    $\mathbb{P}({\bm{g}}\in\Omega,{\bm{x}}_{0}\in\Xi)=\int_{\Omega}\int_{\Xi}\bar{J}({\bm{g}};{\bm{x}}_{0}){\rm d}v_{\bar{\mathcal{J}}}({\bm{g}}){\rm d}{\bm{x}}_{0};$ (23)
@@
    $\bar{J}({\bm{g}};{\bm{x}}_{0})\propto\frac{1}{J({\bm{g}},{\bm{x}}_{0})},{\bm{g}}\in\mathcal{G},{\bm{u}}(k)={\bm{g}}({\bm{x}}(k)),{\bm{x}}(0)\!=\!{\bm{x}}_{0},$ (24)
    where $J({\bm{g}},{\bm{x}}_{0})$ denotes the cost function evaluated with control law $\bm{g}$} and initial state ${\bm{x}}_{0}$.
@@
    Boundedness: $v_{\bar{\mathcal{J}}}(\bar{\mathcal{J}})=\bar{M}_{\bar{\mathcal{J}}},~\bar{J}_{\rm min}\leq\bar{J}({\bm{g}};{\bm{x}}_{0})\leq\bar{J}_{\rm max}.$
Here we note that, through the aforementioned five properties, we establish a space $\bar{\mathcal{J}}$, which serves as a tool for indirectly learning the unknown cost function $J(\cdot)$. Although there are various possibilities for the type of the unknown cost function, under the boundedness assumption 2, we can construct a function $\bar{J}({\bm{g}};{\bm{x}}_{0})$ belonging to the space $\bar{\mathcal{J}}$ and related to the cost function $J(\cdot)$. The inverse in (24) is used to convert the cost function into a reward function, which allows higher reward regions to have higher sampling probability according to TS mechanism. This facilitates the following learning mechanism design and learning convergence analysis.
@@
$\bar{J}^{t}:=\mathbb{E}\left[\bar{J}\right]:=\int_{\bar{\mathcal{J}}}\bar{J}F_{\rm pdf}^{t}({\rm d}v_{\mathcal{J}}(\bar{J})),$ (29)
and the restricted and normalized predictive density at the $t$-th segment can be obtained as
$\bar{J}^{t}_{\Omega}:=\int_{\Omega}\bar{J}\Pi_{\Omega}^{t}({\rm d}v_{\mathcal{J}}(\bar{J})),$ (30)
where $\Pi_{\Omega}^{t}$ is the posterior distribution restricted and normalized to the set $\Omega$, given by
$\Pi_{\Omega}^{t}(\cdot)=\frac{\int_{\cdot\cap\Omega}R_{t}F_{\rm pdf}^{0}({\rm d}v_{\mathcal{J}}(\bar{J}))}{\int_{\Omega}R_{t}F_{\rm pdf}^{0}({\rm d}v_{\mathcal{J}}(\bar{J}))}.$ (31)
@@
Based on the distribution function $F_{\rm pdf}^{t}(\bar{J})$, the detailed procedure of the TS adaptive segment-length learning algorithm is illustrated in Algorithm 2. The algorithm first sets the prior distribution, samples the performance function, and selects the initial control law. When no other prior knowledge is available, the prior distribution can be chosen as a uniform distribution.
@@ IV-C Convergence Analysis of the Learning Process
$T_{d}(x)$ $:=\left\{\!\!\begin{array}[]{ll}\log x,&\text{if}~d=D_{\rm KL},\\ \sqrt{x}-1,&\text{if}~d=\frac{1}{2}D^{2}_{\rm H}.\end{array}\right$. (35)
@@
$\exists\epsilon_{L},P_{0},\delta_{\bar{J}}\!:\!\mathbb{P}(\bar{J}^{T}({\bm{g}})\in{\mathcal{B}}^{c}(\bar{J}^{*},\delta_{\bar{J}}))\!<\!P_{0}\exp(-T\epsilon_{L}),\text{a.s.},$ (37)
@@
$\lim\limits_{t\rightarrow\infty}\inf d(\bar{J}^{t}_{\Omega},\bar{J}^{*})>0,~\text{where}~\Omega=\mathcal{B}^{\rm c}(\bar{J}^{*},\delta_{\bar{J}});$ (38)
@@
The proof of Theorem 2 is divided into two steps. In the first step, a martingale process is constructed, and several properties of the martingle are analyzed. In the second step, the proof of Theorem 2 is completed by using the properties of the constructed martingale process. In the following analysis, the complementary set of the neighborhood ${\mathcal{B}}(\bar{J}^{*},\delta_{\bar{J}})$ is denoted as $\Omega={\mathcal{B}}^{c}(\bar{J}^{*},\delta_{\bar{J}})$.
@@
Specifically, from the definition of Hellinger distance (6) and KL divergence (7), the term $d(\bar{J}^{t}(\cdot),\bar{J}^{*}(\cdot))$ is bounded. Condition (39) ensures that an infinite $T_{d}^{t}$ does not exist, i.e., $T_{d}^{t}$ is also bounded. Thus the inequality (48) holds since the sum of multiple bounded items is also bounded. Moreover, according to (47), we have
$M_{T+1}=M_{T}+T_{d}^{T+1}+d(\bar{J}^{T+1}(\cdot),\bar{J}^{*}(\cdot)).$
Using (46), the expectation of $M_{T+1}$ can be written as
$\mathbb{E}[M_{T+1}]$ $=M_{T}+\mathbb{E}[T_{d}^{T+1}+d(\bar{J}^{T+1}(\cdot),\bar{J}^{*}(\cdot))]$
$=M_{T}+0=M_{T},$
which leads to the equation (49).
@@
$\lim\limits_{T\rightarrow\infty}\frac{1}{T}\sum\limits_{t=1}^{T}\!\left[T_{d}\left(\!\frac{L_{t}(\Omega)}{L_{t-1}(\Omega)}\!\right)\!+\!d(\bar{J}^{t}(\cdot),\bar{J}^{*}(\cdot))\!\right]\!=\!0,\text{a.s.}$ (51)
@@
$\exists\epsilon_{L}(\Omega):~L_{T}(\Omega)<L_{0}(\Omega)\exp(-T\epsilon_{L}),\text{a.s.}$ (53)
@@
$\Rightarrow L_{T}(\Omega)$ $<L_{0}(\Omega)\exp(-T\epsilon_{L}),$ (57)
@@
$\lim\limits_{t\rightarrow\infty}\mathbb{P}(\bar{J}\in\Omega)=0,~\text{a.s.}$ (65)
@@
The first condition (38) in Theorem 2 requires that outside the neighborhood of $\bar{J}^{*}$ (specifically, $\Omega(\delta_{\bar{J}})=\mathcal{B}^{\mathrm{c}}(\bar{J}^{*},\delta_{\bar{J}})$), the distance $d(\bar{J},\bar{J}^{*})$ must always be lower bounded by some positive constant. Intuitively, this is essentially an “identifiability” or “positive information gain” condition: as long as the predictive density $\bar{J}$ is discernibly different from the true density $\bar{J}^{*}$ (i.e., lies outside the $\delta_{\bar{J}}$-ball), new data can clearly distinguish the current controller from the optimal controller, meaning that each learning step yields strictly positive information gain. In practical control applications, this condition is generally satisfied, because different controllers usually lead to different closed-loop state-input trajectories, which in turn result in distinguishable long-term cost distributions. Moreover, in Theorem 2, $\delta_{\bar{J}}$ is a problem-dependent constant rather than a tunable hyperparameter. For a given system and reward function $\bar{J}(\cdot)$, $\delta_{\bar{J}}$ is fixed by the condition (38).
The second condition (39) requires that the variance sequence $\{\mathbb{V}(T_{d}^{t})\}$ is square-summable, that is, $\sum_{t=1}^{\infty}t^{-2}\mathbb{V}(T_{d}^{t})<\infty$. Since each learning segment has a bounded length and the system state, input, and disturbance are all assumed to have bounded second moments (i.e., the working region is bounded and the noise variance is finite), both the cost in each segment and the variance of $T_{d}^{t}$ are uniformly bounded by some constant. Essentially, this is a “light-tailed noise” requirement on the observation noise, which is met in most practical closed-loop experimental scenarios.
In Theorem 2, $\delta_{\bar{J}}$ influences both the neighborhood size and the convergence rate. By definition (8), $\delta_{\bar{J}}$ is the radius of the neighborhood $\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}})$, so the measure $v_{\bar{\mathcal{J}}}(\mathcal{B}(\bar{J}^{*},\delta_{\bar{J}}))$ increases monotonically with $\delta_{\bar{J}}$: a larger $\delta_{\bar{J}}$ means a larger neighborhood and thus worse pointwise accuracy. The convergence rate $\epsilon_{L}$ of the learning process (37) is determined by the dynamics of the likelihood mass on the complement $\Omega=\mathcal{B}^{\rm c}(\bar{J}^{*},\delta_{\bar{J}})$. In general, a larger $\delta_{\bar{J}}$ corresponds to a smaller $\Omega$. Typically, when the current estimate lies in $\Omega$, the empirical reward tends to be smaller relative to $\bar{J}^{*}$ (as shown in (42)), which leads to a smaller likelihood ratio in the convergence-rate expression and thus a larger $\epsilon_{L}$. Therefore, a larger $\delta_{\bar{J}}$ typically results in a larger $\epsilon_{L}$, namely, a faster exponential convergence rate.
According to Theorem 2, it can be claimed that after a finite number of updates, the predictive density $\bar{J}^{t}$ converges to a neighborhood of the true reward $\bar{J}^{*}$ with high probability. This is analyzed in the following corollary.
Corollary 2.
For a constant $P_{\rm fail}\in(0,1)$, if the number of segments $T$ satisfies
$T>\frac{1}{\epsilon_{L}}\log\frac{P_{0}}{P_{\rm fail}},$ (68)
then the probability that the predictive density $\bar{J}^{T}$ converges to a neighborhood of the true reward $\bar{J}^{*}$ satisfies
$\mathbb{P}(\bar{J}^{T}\in{\mathcal{B}}(\bar{J}^{*},\delta_{\bar{J}}))\geq 1-P_{\rm fail}.$ (69)
Proof.
For $T>\frac{1}{\epsilon_{L}}\log\frac{P_{0}}{P_{\rm fail}}$, we have
$P_{0}\exp(-T\epsilon_{L})<P_{\rm fail}.$ (70)
Then according to Theorem 2, we obtain
$\mathbb{P}(\bar{J}^{T}\notin{\mathcal{B}}^{\rm c}(\bar{J}^{*},\delta_{\bar{J}}))<P_{0}\exp(-T\epsilon_{L})<P_{\rm fail},$ (71)
based on which the conclusion (69) can be proved.